### v1

# GAGS: Granularity-Aware Feature Distillation for Language Gaussian Splatting

Abstract
3D open-vocabulary scene understanding, which accurately perceives complex semantic properties of objects in space, has gained significant attention in recent years. In this paper, we propose GAGS, a framework that distills 2D CLIP features into 3D Gaussian splatting, enabling open-vocabulary queries for renderings on arbitrary viewpoints. The main challenge of distilling 2D features for 3D fields lies in the multiview inconsistency of extracted 2D features, which provides unstable supervision for the 3D feature field. GAGS addresses this challenge with two novel strategies. First, GAGS associates the prompt point density of SAM with the camera distances, which significantly improves the multiview consistency of segmentation results. Second, GAGS further decodes a granularity factor to guide the distillation process and this granularity factor can be learned in a unsupervised manner to only select the multiview consistent 2D features in the distillation process. Experimental results on two datasets demonstrate significant performance and stability improvements of GAGS in visual grounding and semantic segmentation, with an inference speed 2$\times$ faster than baseline methods. The code and additional results are available at https://pz0826.github.io/GAGS-Webpage/.
Figure 1: Given multiview images of a scene, GAGS learns a 3D Gaussian field associated with semantic features, which enables accurate open-vocabulary 3D visual grounding in the scene.
1 Introduction
3D scene understanding is a fundamental task in computer vision and a critical challenge in fields like robotics [15, 35, 38] and autonomous driving  [17, 45]. Recent advances in artificial intelligence and deep learning have driven research on open-vocabulary scene understanding, enabling users to query scene models using natural language. This enhances the efficiency and intuitiveness of human-computer interaction, fostering closer integration between intelligent systems and human cognitive processes.
Figure 2: A 3D point shows different semantics at different scales.
Due to the lack of large-scale, diverse 3D scene datasets with language annotations, current efforts focus on extending the knowledge of 2D vision-language models to 3D scenes. Early methods, such as OpenScene [29], compute pixel-level embeddings using pre-trained segmentation models and project them onto 3D point clouds. Recent approaches like LERF [19] employ NeRF [27] to represent 3D scenes, directly integrating CLIP [32] features into the scene modeling. Another recent work LangSplat [30] extends the open-vocabulary feature learning to the 3D Gaussian Splatting (3DGS) [18] technique, which enables fast rendering and impressive performances of visual grounding with arbitrary text descriptions.
A noticeable challenge of learning such open-vocabulary features for 3D fields lies in the multiview inconsistency of the extracted 2D features. As shown in  Fig. 2, given a specific 3D location, it can be described as “desktop”, “bowl”, and “Ramen” at the same time according to the different granularities. In this case, the extracted 2D features show no multiview consistency on different viewpoints. Brute-force training a 3D feature field from inconsistent 2D features leads to averaged and degenerated 3D features. LangSplat [30] resolves this problem by learning three different feature fields corresponding to different segmentation granularity of SAM [20]. However, learning three different fields increases both training memory and time cost and requires the query texts to be compared with three different features. N2F2 [2] proposes to merge the features of different granularity by selecting the most activated features within a set of manually predefined descriptions. In this case, the performance is strongly related to the quality of these predefined descriptions. If the granularity of query descriptions differs from these predefined descriptions, the performance degenerates severely.
Figure 3: In comparison to SAM with default prompt points, we associate the prompt point density with distance to the object, which greatly enhances the consistency of multi-view segmentation.
In this work, we introduce Granularity-Aware 3D Feature Learning for Gaussian Splatting (GAGS), a framework built on 3DGS that incorporates granularity-aware CLIP features in the reconstructed 3D Gaussian field. GAGS essentially consists of two strategies to overcome the inconsistency of multiview 2D features.
First, we improve the segmentation granularity to improve the consistency. In the CLIP feature extraction, we need to apply the SAM to segment images into a region and extract a CLIP feature for the zoomed segmented region. SAM may have multiple possible solutions to segment an image, which introduces inconsistency for segmentation from different viewpoints. As shown in  Fig. 3, we improve the consistency of SAM segmentation results among multiview images by introducing a granularity-aware prompting strategy. SAM relies on the prompt points to determine the segmentation granularity. To keep the multiview consistency, images with far-away viewpoints should have dense prompt points while images with nearby viewpoints should have sparse prompt points. Thus, we associate the prompt point density with the camera distance to the subject by utilizing the reconstructed 3D GS. We found that this adaptive sampling effectively results in consistent segmentation results on multiview images and improves the multiview consistency of CLIP features.
Figure 4: Our granularity-select feature learning. Our method leverages the inherent consistency of 3D Gaussian splatting to perform granularity-aware feature distillation, enhancing the stability and accuracy of learned object features.
Second, we incorporate a granularity decoder on the learned 3D feature to decode a granularity factor as shown in  Fig. 4, which helps the distillation process only learn the multiview consistent 2D features while neglecting inconsistent ones. On every image, we simultaneously extract three levels of 2D CLIP features, i.e. sub-parts, parts, and objects. Then, the decoded granularity factor will be incorporated in the distillation loss to select the CLIP feature of a specific granularity for training. Note that the granularity scale is automatically learned in the distillation process without additional supervision.
We conduct experiments on the LERF [19] dataset and a self-annotated Mip-NeRF-360 dataset for evaluation. Experimental results demonstrate that our method outperforms baseline methods in the open-vocabulary localization and semantic segmentation tasks with a 2$\times$ improvement in query speed.
2 Related work
2.1 Point-based Radiance Field
Radiance fields have long been used to represent 3D scenes [12, 22]. Recently, neural radiance field methods, exemplified by NeRF [27], have achieved breakthroughs in rendering quality, bringing renewed attention to this field. Trained NeRF models leverage neural networks to learn complete scene representation, enabling high-quality novel view rendering. However, due to the reliance on ray sampling and deep MLP architectures, NeRF suffers from slow training and rendering speeds. While various approaches [28, 11] have attempted to improve these aspects, they still struggle to balance speed and quality. Additionally, NeRF’s implicit representation makes it difficult to directly obtain the scene geometry. Methods such as NeuS [39] and VolSDF [41] address this by introducing Signed Distance Functions (SDF) to represent geometric surfaces, deploying new volumetric rendering techniques to learn the SDF representation. 3D Gaussian Splatting [18] introduces a novel solution as a point-based radiance field method, representing scenes using discrete 3D Gaussian distributions. It replaces expensive random-sampling-based volumetric rendering with $\alpha$-blending, significantly boosting rendering speed. Additionally, the explicitly stored Gaussian kernels allow direct retrieval of scene geometry after training. Leveraging these advantages, we adopt 3DGS as the framework and extend it to support open-vocabulary 3D scene understanding.
2.2 SAM
As one of the most impressive foundational vision models, Meta’s SAM [20] has demonstrated exceptional zero-shot 2D segmentation capabilities. SAM supports flexible prompts, allowing it to generate multi-level masks for target objects based on inputs such as points, bounding boxes, masks, and text. Additionally, SAM offers the capability to automatically segment the entire image using predefined dense grid point prompts. Numerous methods have since emerged to extend SAM’s capabilities to 3D space. Anything-3D [34] elevates SAM’s segmentation to 3D, while Part123 [24] uses SAM’s segmentation masks to reconstruct 3D models with high-quality segmented parts. Feature 3DGS [46] and Gaussian Grouping [42] integrate SAM’s features and object segmentation results into 3D Gaussians, achieving high-quality segmentation of novel views and 3D scenes. Semantic Gaussian [13] utilizes various prompts to obtain instance-level segmentations for feature extraction. We further optimize SAM’s segmentation process by incorporating scene geometry, guiding it to generate masks with stronger multi-view consistency, resulting in more accurate features for training 3D language fields.
2.3 Open-vocabulary Scene Understanding
Scene understanding is a basic task in the field of computer vision. With the rapid advancement of deep learning, numerous methods [40, 4, 5, 37] have made significant progress across various subtasks of scene understanding. However, the limited availability of 3D training data has long posed a challenge for achieving comprehensive 3D scene understanding. Vision foundation models like CLIP [32] have opened new avenues for open-vocabulary scene understanding. Due to the image-aligned nature of CLIP features, subsequent works such as CLIP2Scene [6] and Openscene [29] directly leverage CLIP-based 2D scene understanding models [9, 23] to obtain dense features with CLIP semantics. Approaches like [44, 25, 19, 47, 36] generate dense semantic features by performing multi-level image cropping and feature fusion, while often incorporating pixel-aligned feature supervision, such as DINO [3], to address the blurriness in semantic boundary. Some other researches [21, 14, 16, 30] utilize pre-trained image segmentation model [23, 7] to obtain semantically meaningful object-level patches, enabling more accurate scene understanding.
Figure 5: Pipeline of our method. Given a set of images with camera poses, our method first uses 3D Gaussian Splatting to reconstruct the scene’s geometric representation, then utilizes it for granularity-aware segmentation and CLIP feature distillation. The finally output 3D feature field supports open-vocabulary queries.
3 Method
Given multiple input images with known camera poses, our target is to learn a 3D feature field represented by 3D Gaussians with associated feature vectors. Then, we can render arbitrary novel-view images from this 3D feature field and also use text prompts to query the 3D feature field for a range of downstream tasks, such as object localization or semantic segmentation.
An overview of our method is provided in Fig. 5. Inspired by LangSplat [30], We first train a 3D Gaussian field from the input posed multiview images. Then, for every input image, we apply the SAM to get a set of segmented regions. Each region is cropped and fed into the CLIP image encoder to get a CLIP feature, which is regarded as the feature for all pixels in the segmented region. Finally, we associate a trainable feature vector on each Gaussian and apply the splatting technique to render a feature map from the 3D Gaussian field. By minimizing the difference between the rendered feature map and the extracted CLIP feature map, we train these associated feature vectors on Gaussians. The main challenge lies in extracting and learning multiview-consistent CLIP features. In GAGS, we propose the following Granularity-aware Segmentation and Granularity-aware Distillation to address this challenge.
3.1 Granularity-aware Segmentation
The extracted CLIP features are strongly related to the region size of the segmentation results produced by SAM, different segmentation granularity leads to totally different CLIP features. However, since the SAM is applied to every input image separately and the same object may show different sizes on different viewpoints, the segmentation results of SAM are already multiview inconsistent, resulting in inconsistent CLIP features. For example, the ramen bowl in Fig. 3(c)(d) corresponds to different segmentation granularities across various viewpoints. We propose a granularity-aware segmentation to improve the multiview consistency of segmentation results.
Observations. The segmentation granularity of SAM is controlled by the prompt point density. In the ramen bowl region of Fig. 3(d)(f), Dense prompt points usually lead to small segmentation regions while sparse prompt points result in large segmentation regions. Another observation is that if the object is far away from the camera, the object is small on the image and the segmentation mask should be small while nearby objects should be large on the image. In this case, the prompt point density is depth-dependent, where we should use a higher prompt point density for distant viewpoints to segment small objects while a lower density for nearby viewpoints to segment the same large objects.
Figure 6: Our granularity-aware segmentation module utilizes the Gaussian kernel point cloud and camera poses to guide SAM and CLIP in producing multi-view consistent semantic features.
Granularity-aware prompt point density. We compute the prompt point density from the depth maps as follows. As illustrated in Fig. 6, Given an input image, we begin by dividing it into a grid of patches and our target is to determine how many sample points should lie in each patch. Since the rendered depth determines the density, we first render a depth map $D$ from the 3D Gaussian field. Then, we convert the depth map into a set of depth points and reproject these depth points onto all the other visible views. Then, we find the minimum depth values for all the visible views, denoted by $MD$ which maps a pixel to a minimum visible depth value of all visible views. Then, the sample point number for a specific patch $P$ is computed by
$n_{P}=\frac{1}{|P|}\sum_{p\in P}\frac{MD^{2}(p)}{D^{2}(p)}\cdot n,$ (1)
where $n_{P}$ is the number of prompt points in $P$, $|P|$ is the pixel number in the patch, $p\in P$ is a pixel, $MD^{2}(p)$ is the squared minimum depth value of pixel $p$ while $D^{2}(p)$ is the squared depth value of pixel $p$, and $n$ is a predefined prompt point number.
Explanation of Eq. 1. The ratio $\frac{MD^{2}(p)}{D^{2}(p)}\in(0,1]$ controls the prompt point density on a specific view. When this view is the one with the smallest viewing distance, the ratio is 1.0, which means we sample $n$ points on the nearest view. While, for a far-away viewpoint, we use fewer prompt points than the nearest view because the object would be smaller on these far-away views. In this way,  Eq. 1 results in multiview consistent prompting for the SAM model on different viewpoints, which increases the multiview consistency of the segmentation results. Moreover, we find that using random prompt points in the patch produces better results than the even prompt points.
Figure 7: Our granularity-aware language field distillation module introduces optimizable low-dimensional semantic features into the pre-trained Gaussian scene. The rendered semantic features are fed into the granularity decoding head, which selects an appropriate scale for each pixel. These selected features are then used to compute loss with the output full-dim semantics of the feature decoder for optimization.
3.2 Granularity-aware Distillation
Though the above granularity-aware segmentation improves the consistency, the prediction of SAM still exists with randomness and leads to inconsistency. Thus, we further propose a novel granularity-aware distillation to autonomously select the granularity with the best multi-view consistency during optimization. Given the granularity-aware prompt points, SAM returns three levels of segmentation results and on all three levels, we extract CLIP feature maps $f_{s}$, $f_{p}$ and $f_{w}$ corresponding to “sub-part”, “part”, and “whole”. We design a strategy to let the distillation automatically select the most multiview-consistent features.
Decoding granularity factor. Given the 3D Gaussian field associated with trainable feature vectors, we apply the splatting technique to render a feature map for a training viewpoint. We adopt a feature decoder to decode a granularity factor and a predicted CLIP feature from the rendered feature vectors by
$f_{\text{clip}},\eta=\mathcal{D}(f_{\text{render}}),$ (2)
where $f_{\text{clip}}$ is the predicted CLIP feature, $\eta\in\mathbb{R}^{3}$ is the predicted granularity factor, $\mathcal{D}$ means the MLP-based decoder shared among all viewpoints, and $f_{\text{render}}$ is the rendered feature.
Weighted distillation loss. We convert the granularity factor into weight by applying the softmax operator between three scores $w=\text{softmax}(\eta)$, which corresponds to the weight of the “sub-part”, “part”, and “whole” respectively. Then, we will train with a weighted L2 loss for distillation
$\ell_{\text{distill}}=\sum_{n\in\{s,p,w\}}w_{n}\|f_{\text{clip}}-f_{n}\|_{2}^{2},$ (3)
where $n$ means the name, $s,p,w$ correspond to “sub-part”, “part”, and “whole” respectively, and $w_{n}\in\mathbb{R}$ means the weight for the current level. We do not directly supervise $w_{n}$ but let the optimization process automatically select the best weight to learn the granularity of the feature.
Entropy regularization. To further encourage the weight to converge to a single scale instead of learning an averaged feature, we adopt an entropy regularization term,
$\ell_{\text{entropy}}=-\sum_{n\in\{s,p,w\}}w_{n}\log w_{n},$ (4)
which encourages the weight to be either 0 or 1. By adopting the granularity-aware distillation, GAGS only learns one 3D feature field instead of 3 feature fields in LangSplat [30], resulting in a more compact representation. The predicted granularity factor enables adaptively adjusting the granularity for different views and selecting the most multiview-consistent one.
Region-aware weighted distillation. In our experiments, we observed that due to perspective variations and intrinsic object scale differences, different objects occupy varying proportions within the same image. During loss calculation, objects with a larger region in the image contribute more significantly to the loss, thereby dominating the optimization direction of semantic features. This imbalance may hinder feature learning for smaller objects. To address this issue, we propose a region-aware distillation loss, normalizing the loss by the region size of each object to ensure that all objects contribute equally to the total loss
$\ell_{\text{r-distill}}=\alpha_{r}{\ell_{\text{distill}}},\alpha_{r}=\frac{\sum_{i=1}^{n_{r}}{S}({R}_{i})}{n_{r}\cdot{{S}({R})}},$ (5)
where $\alpha_{r}$ denotes the region-aware factor, with $n_{r}$ as the number of objects in the image and ${S}({R})$ as the area of each region.
Feature consistency loss. Inspired by the process of contrastive learning, we further introduce a feature consistency loss to encourage the features inside the same segmentation region to be consistent with each other
$\ell_{\text{cons}}=\sum_{i=1}^{n_{r}}\sum_{p\in R_{i}}\frac{(f_{p}-\overline{f_{R_{i}}})^{2}}{S(R_{i})}$ (6)
where $f_{p}$ and $\overline{f_{R_{i}}}$ represent the $f_{\text{clip}}$ of pixel $p$ and the average $f_{\text{clip}}$ of the $i$-th region. By pulling together features of the same object region across each viewpoint, we ultimately enhance the feature consistency for the 3D Gaussians belonging to the same object.
In summary, the optimization loss $\mathcal{L}$ is
$\mathcal{L}=\ell_{\text{r-distill}}+\lambda_{\text{entropy}}\ell_{\text{entropy}}+\lambda_{\text{cons}}\ell_{\text{cons}}$ (7)
Method LERF Mip-NeRF360
Ramen Teatime Kitchen Figurines Overall Room Counter Garden Bonsai Overall
GS-Grouping 32.39 69.49 50.00 44.64 49.13 79.31 56.76 57.14 66.67 64.97
LEGaussian 69.01 79.66 63.64 57.14 67.36 55.17 72.97 71.42 61.11 65.17
GOI 56.33 67.80 68.18 44.64 59.24 68.97 67.57 66.67 72.22 68.86
Langsplat 63.38 88.14 81.82 76.79 77.53 75.86 91.89 52.38 72.22 73.09
GAGS(Ours) 69.01 88.14 90.91 78.57 81.66 93.10 97.30 80.95 83.33 88.67
Table 1: Quantitative comparisons of 3D object location on the LERF and Mip-NeRF 360 dataset. We report the mean accuracy(%↑).
Method LERF Mip-NeRF360
Ramen Teatime Kitchen Figurines Overall Room Counter Garden Bonsai Overall
LEGaussian 20.17 32.29 22.3 23.41 24.54 25.49 35.26 33.18 22.29 29.06
GS-Grouping 26.39 53.97 31.33 34.55 36.56 54.38 47.67 40.36 54.13 49.14
GOI 33.69 55.79 54.46 23.85 41.95 60.26 46.64 59.75 67.26 58.48
Langsplat 45.82 62.47 43.39 51.10 50.70 51.22 64.69 49.83 52.96 54.68
GAGS(Ours) 46.81 60.29 55.8 53.59 54.12 65.16 61.05 61.22 70.52 64.49
Table 2: Quantitative comparisons of 3D semantic segmentation on the LERF and Mip-NeRF 360 dataset. We report the mean IoU(%↑).
4 Experiments
4.1 Experimental Protocol
Datasets. We use the LERF dataset [19] and a self-annotated Mip-NeRF-360 [1] dataset for evaluation. The LERF [19] dataset comprises 3700+ phone-captured images from 14 different scenes. LangSplat [30] provides text descriptions and corresponding multi-view segmentation masks within four scenes: “ramen”, “waldo_kitchen”, “teatime”, and “figurines”, aiming to evaluate text-based object localization and segmentation. The Mip-NeRF-360 [1] dataset provides multi-view images for 9 indoor and outdoor scenes. Following LangSplat [43], we annotated four of the most complex scenes: “room” , “counters”, “garden” and “bonsai”. For each scene, we provided query texts for multiple types of objects along with the corresponding multi-view segmentation masks.
Baselines. We adopt recent relevant works including GS-Grouping [42], LEGaussian [36], GOI [31], and LangSplat [30] as our baseline methods. We employ GSplat [43] to construct initial Gaussian fields and evaluate all methods in the same setting for fair comparisons.
Metrics. We follow LERF [19] and LangSplat [30] to evaluate text-based 3D localization and segmentation accuracy on multi-view images. For localization, we evaluate the mean accuracy (mAcc) of the predicted locations falling within the ground truth bounding boxes. For segmentation, we assess the mean Intersection over Union (mIoU) between the predicted and the ground truth masks.
Figure 8: The relevance visualization results for open-vocabulary queries. Each row from top to bottom represents Ground Truth, Gaussian Grouping, GOI, Langsplat, and our method. Below each column is the corresponding input text description.
Implementation Details. We utilized the SAM ViT-H [10] and OpenCLIP ViT-B/16 [8] for segmentation and CLIP feature extraction across the multi-view images. All experiments are conducted on a single RTX-4090 GPU. More details on implementation are provided in the supplementary material.
4.2 Comparison with Baseline Methods
In  Fig. 8, we visualize the segmentation results of the baseline methods and our method. Tab. 1 and Tab. 2 report the localization and segmentation accuracy of these methods on the LERF and Mip-NeRF-360 datasets. All performances reported were re-evaluated and the metrics of Langsplat has some differences from the original paper, we provide further analysis in the supplementary material.
As shown in  Fig. 8, GS Grouping [42] achieves sharp segmentation results by fusing multi-view SAM segments. However, segmentation masks for the same object across multiple views might lack consistency, failing to support and merge with each other. The incorrect fusion of segmentation masks can lead to detection failures, which is especially noticeable on small objects. GOI [31] uses detection results from the Grounding-SAM [26] to fine-tune a semantic-space hyperplane for selecting the final segment. However, Grounding-SAM also tends to fail with small and long-tail objects, causing errors in hyperplane optimization and leading to missed detections. LangSplat [30] embeds multi-level CLIP features within its Gaussian field, achieving better localization accuracy by avoiding hard fusion. However, the multi-view CLIP features often lack consistency, which introduces substantial noise into the feature field, yielding erroneous or noisy segmentation outcomes.
In contrast, our method encourages the distillation of multi-view consistent CLIP features into the Gaussian field, avoiding noise from conflicting multi-view features, which leads to a 4.1% and 3.4% improvement in localization and segmentation accuracy on the LERF dataset. Moreover, the proposed Granularity-aware Distillation can adaptively fuse distill multi-granularity features, ensuring the effectiveness of GAGS in multi-scale object localization. Consequently, on the Mip-NeRF-360 dataset, which includes objects of varying scales, we achieved a more significant improvement over the baseline, with a 15.5% mAcc improvement for localization and a 6.0% increase in mIoU for segmentation.
Runtime analysis. In Tab. 3, we report the runtime of baselines and GAGS. Although GOI [31] achieved the shortest training time, it requires fine-tuning a high-dimensional hyperplane during testing to segment the target, resulting in an inference time nearly 100 times that of GAGS. When comparing with LangSplat [30], GAGS achieves a similar training speed. However, LangSplat requires rendering multi-granularity feature maps to compare and segment the target, while GAGS adaptively fuses multi-granularity features, necessitating only a single feature map rendering and comparison. Thus, GAGS is two times faster than LangSplat in inference.
Figure 9: Qualitative results of the ablation study. Below each column is the corresponding input text description.
Method Training (min) Inference (s)
Prep. GS Lang. Total IE Rend. Pred. OSH Total
GS-Grouping 20 10 40 70 66 ~1 ~1 - 68
GOI 10 10 10 30 - ~1 ~1 2342 2344
Langsplat 50 10 30 90 - 15 34 - 49
GAGS(Ours) 25 10 50 85 - 13 11 - 24
Table 3: Time evaluation for each step in training and prediction on the scene ”ramen” of the LERF dataset. During training, “Prep.” refers to the preprocessing stage, while “GS” and “Lang.” represent the training stage of RGB and feature Gaussian fields, respectively. During inference, “IE” denotes obtaining the Identity Encoding of the target, while “Rend.” and “Pred.” refer to rendering feature image and predicting target mask.
4.3 Ablation Studies
In Tab. 4 and Fig. 9, we conduct both qualitative and quantitative ablation evaluations of GAGS on the Mip-NeRF-360 dataset. Model 1 of Tab. 4 is the baseline model, where we use uniform prompt points for SAM segmentation. We then average the CLIP features across different SAM granularities, calculated as $(f_{s}+f_{p}+f_{w})/3$, and distilled the normalized feature into the Gaussian field. Model 2 incorporate the proposed Granularity-aware Distillation (GaD), which boosts the localization and segmentation accuracy by 7.2% and 9.4%, respectively. Further incorporating the proposed Granularity-aware Segmentation (GaS) in Model 3 brings an additional 3.0% mAcc improvement for localization and 2.4% mIoU improvement for segmentation. The results in Fig. 9(a) and (c) also shows that GaS can mitigate noise issues from over-segmentation of nearby objects and low-texture areas, as well as under-segmentation of distant objects.
Moreover, in Models 4-6, we present results by distilling CLIP features using single SAM segmentation granularity. Among them, the part granularity yields the best results, while the subpart granularity performed the worst because the over-segmentation makes CLIP feature extraction unreliable and leads to multi-view conflicts. Model 7, which simply averages the CLIP features from all three SAM granularities, did not yield improvement. In contrast, Model 3, using GaD, can adaptively choose reliable and multi-view consistent features, achieving optimal performance. The comparison between panels (b) and (c) in Fig. 9 also demonstrates that GaD significantly improves segmentation accuracy by distilling multi-view consistent features at the appropriate granularity for each object.
Setting Performance
ID GaS Distill. mIoU(%) mAcc(%)
1 Avg. 53.72 76.84
2 GaD 60.99 86.27
3 GaD 63.94 88.67
4 $f_{s}$ 47.78 77.32
5 $f_{p}$ 61.68 85.74
6 $f_{w}$ 58.63 78.51
7 Avg. 59.45 84.88
Table 4: Ablation Studies on the Mip-NeRF-360 dataset. “GaS” means the proposed granularity-aware segmentation and “Distill.” denotes the feature distillation strategy. For “Distill.”, in addition to distilling the averaged features across all granularities, we also conducted experiments for each individual granularity feature.
5 Conclusion
In this paper, we propose GAGS, a granularity-aware 3D Gaussian Splatting framework for efficient and accurate open-vocabulary queries. By introducing a granularity-aware segmentation and feature distillation strategy during the learning of 3D feature field, GAGS is able to learn clear and multiview consistent semantic features. It avoids the high query cost of rendering multiple language fields and mitigates the impact of outliers. Experimental results across several datasets demonstrate that our method outperforms existing approaches in both performance and speed.